# HybridFlow: Resource-Adaptive Subtask Routing for Efficient Edge-Cloud LLM Inference

### v3

@@ Abstract
Edge-cloud collaborative inference is becoming a practical necessity for LLM-powered edge devices: on-device models often cannot afford the required reasoning capability, while cloud-only inference could be prohibitively costly and slow under strict latency and token/API budgets. However, existing edge-cloud collaboration methods often route per query or fixed steps simply based-on the estimated difficulty. Such coarse and static heuristics overlook subtask dependencies, missing opportunities for parallel execution and budget-adaptive routing. To this end, we propose HybridFlow, a resource-adaptive edge-cloud inference framework that (i) builds a dependency-aware DAG for each query and executes newly unlocked subtasks in parallel, reducing end-to-end latency; (ii) routes each subtask online to the edge or cloud via a learned benefit–cost utility model that dynamically trades accuracy gains against token/API and latency budgets, thereby reducing unnecessary cloud usage while preserving reasoning quality. Across GPQA, MMLU-Pro, AIME24, and LiveBench-Reasoning, HybridFlow improves the cost-accuracy trade-off, reducing latency and cloud API usage while maintaining competitive accuracy against strong structured reasoning baselines. Code: HybridFlow
@@ 2 Related Work
Algorithm 1 HybridFlow: Dependency-Aware Scheduling with Budget-Adaptive Routing
Input: Query $Q$; edge model $M_{\mathrm{edge}}$; cloud model $M_{\mathrm{cloud}}$; planner $M_{P}$; budgets $C_{\max}$; router $f_{\theta}$; step size $\eta$; threshold params $(\tau_{0},\gamma)$
@@
Stage 1: Decompose and initialize
$(T,E)\leftarrow\mathrm{Decompose}(Q;M_{P})$
; build DAG $G=(T,E)$
Initialize in-degrees $\deg(\cdot)$ on $G$; answers map $A[\cdot]\leftarrow\emptyset$
Initialize ready queue $\mathcal{F}\leftarrow\{t\in T:\deg(t)=0\}$
Initialize cumulative cost $C_{\mathrm{used}}\leftarrow 0$ and dual variable $\lambda\leftarrow 0$
Stage 2: Schedule, route, and execute
while
$\mathcal{F}$ not empty do
   Pop a ready subtask $t$ from $\mathcal{F}$
$x\leftarrow\big(t,\{A[p]\}_{p\in\mathrm{Dep}(t)}\big)$
$\hat{u}\leftarrow\sigma(f_{\theta}(\mathrm{Embed}(t)))$ ;
@@
 // Eq. ($0,1)\tau\leftarrow\mathrm{clip}(\tau_{0}+\gamma\lambda,$0,1) ;
 // Eq. (11)if $\bar{u}>\tau$ and $C_{\mathrm{used}}+c(t)\leq C_{\max}$ then
    $A[t]\leftarrow M_{\mathrm{cloud}}(x)$;  $C_{\mathrm{used}}\leftarrow C_{\mathrm{used}}+c(t)$
@@
    $A[t]\leftarrow M_{\mathrm{edge}}(x)$
 foreach $s\in\mathrm{Succ}(t)$ do
    $\deg(s)\leftarrow\deg(s)-1$; if $\deg(s)=0$ then
        push $s$ into $\mathcal{F}$
Stage 3: Aggregate
$R\leftarrow\mathrm{Aggregate}(\{A[t]\}_{t\in T}\ \text{in topological order})$
return
$R$
@@ 3.1 Preliminaries
  • $\Delta k_{i}$: additional API usage cost (in tokens or price units) incurred by cloud execution.
In practice, $\Delta q_{i}$ is estimated offline using automatic metrics such as BartScore (Yuan et al., 2021) by comparing cloud and edge outputs against references or high-quality pseudo-labels. This profiling is done once and does not affect online latency.
To place latency and API usage on a common scale, we define a normalized offloading cost.
@@
For each subtask $t_{i}$, the normalized cost of offloading is
@@
where $l_{\max}^{\mathrm{sub}}$ and $k_{\max}^{\mathrm{sub}}$ are per-subtask upper bounds on additional latency and API cost. Typically $\Delta l_{i}\leq l_{\max}^{\mathrm{sub}}$ and $\Delta k_{i}\leq k_{\max}^{\mathrm{sub}}$, so $c_{i}$ is bounded and comparable across subtasks.
We use a binary variable $r_{i}\in\{0,1\}$ to indicate whether $t_{i}$ is offloaded to the cloud ($r_{i}=1$) or executed on the edge ($r_{i}=0$). Intuitively, subtasks with larger $\Delta q_{i}$ and smaller $c_{i}$ are more attractive to offload.
@@ 3.3 Utility-based Subtask Routing
$\hat{u}_{i}=\sigma\!\big(f_{\theta}(z_{i})\big)\in(0,1),$ (8)
where $\sigma$ is the sigmoid function. The estimate $\hat{u}_{i}$ is trained to approximate the ideal utility $u_{i}$ in Def. 3.2.
@@
where $\eta>0$ is a step size and $C_{\mathrm{used}}(t)=\sum_{j\leq t}r_{j}c_{j}$ is the cumulative normalized cost used so far. When the system overspends, $\lambda_{t}$ increases and the routing policy becomes more conservative.
@@ 4.1 Experiment Setup
We evaluate HybridFlow on four reasoning benchmarks spanning mathematical, scientific, and general domains: GPQA (Rein et al., 2024); AIME24; MMLU-Pro (Wang et al., 2024b), and LiveBench-Reasoning (White et al., 2025). We select three metrics that jointly capture reasoning quality and system efficiency. (i) $Acc$ measures the correctness of reasoning by comparing model outputs with gold answers across all benchmarks. (ii) $C_{\mathrm{time}}$ denotes the end-to-end inference latency per query, including decomposition, routing, and execution. (iii) $C_{\mathrm{API}}$ quantifies the number of tokens consumed by cloud LLM calls, reflecting token efficiency and cost.
@@
HybridFlow deploys Llama3.2-3B on the edge device in two roles: (i) as the planner $M_{P}$ that decomposes each query into a task-level DAG, and (ii) as the edge executor $M_{\mathrm{edge}}$ that runs subtasks not offloaded. Each subtask is encoded with qwen3-embedding-0.6b into an embedding $z_{i}$, which is fed to a lightweight router: a two-hidden-layer MLP that predicts an offline utility estimate $\hat{u}_{i}$. The router is offline warm-started with AdamW (learning rate $1\times 10^{-4}$) by regressing to profiled utility targets. During inference, HybridFlow performs online adaptation with dual-threshold updates that track budget consumption; an online calibration head is updated from partial feedback via a contextual-bandit strategy. Routing decisions use the current threshold to assign each ready subtask to either $M_{\mathrm{edge}}$ or the cloud model $M_{\mathrm{cloud}}$ (GPT-4.1, via API). All edge-side computation (planning, local execution, and embedding) runs on a single NVIDIA RTX 3090 GPU (24GB). All LLM use a fixed temperature of $0.6$.
@@ 4.2 Results
While HybridFlow’s accuracy approaches the top-performing CoT with GPT-4.1, it does so with dramatically lower token consumption and latency. Table 2 shows that HybridFlow achieves the lowest average $C_{\mathrm{API}}$ (0.0088) among all collaborative baselines, indicating high token efficiency in cloud usage. When considering both accuracy and efficiency jointly, HybridFlow consistently dominates alternative hybrid systems: it improves accuracy relative to HybridLLM and DoT while simultaneously reducing latency and API cost. This demonstrates that HybridFlow effectively balances the strengths of edge and cloud models through adaptive routing and parallelized task execution, achieving a superior overall accuracy–efficiency frontier.
Table 3: Ablation of routing strategies on GPQA. We compare offload rate, accuracy, latency, API cost, normalized total cost, and utility (see Sec. 3 for definitions). HybridFlow achieves the best trade-off between accuracy and cost, the highest utility score.
@@ 4.3 Ablation Study
We evaluate our adaptive router against deterministic and naive allocation strategies (Table 3). Using only the edge model (Edge) incurs zero API cost but yields the lowest accuracy (25.54%), whereas executing all subtasks in the cloud (Cloud) achieves the highest accuracy (57.28%) with higher latency and API cost ($C_{\text{API}}{=}0.0185$). The Random baseline improves accuracy (46.00%) by occasionally offloading, but lacks a principled mechanism to prioritize high-value subtasks, resulting in lower utility. In contrast, HybridFlow attains 53.33% accuracy at a much lower API cost ($C_{\text{API}}{=}0.0075$) and achieves the highest utility score (0.7940), indicating a more favorable accuracy–cost trade-off under our unified metric. Overall, these results suggest that the learned router better aligns subtask-level benefit with resource expenditure.
@@ Appendix A Notations
$K_{\max}^{\mathrm{global}}$ Global API usage budget over the entire query.
$L_{\max}^{\mathrm{global}}$ Global latency budget over the entire query.
$k_{\mathrm{used}}$ Cumulative API usage consumed so far during inference.
$l_{\mathrm{used}}$ Cumulative latency incurred so far during inference.
@@ Appendix C Implementation Details
    (Well-formed outputs) At least one node is labeled Generate, and every Generate node has no outgoing edges (i.e., it is a sink in $G(Q)$).
@@
To train the proposed resource-aware router, we construct an offline profiling dataset from 3,0003{,}000 sampled queries drawn from two benchmarks: MMLU-Pro (different from the main test samples) and Math500 (covering general knowledge reasoning, targeting structured, multi-step reasoning). For each query, we perform the complete pipeline of task decomposition, subtask allocation, and execution using both the edge and cloud models. During this process, we record the response quality, inference latency, and API cost for each subtask.
@@
For every subtask $t_{i}$, we estimate the accuracy gap between the edge and cloud models using BartScore (Yuan et al., 2021), which provides a differentiable approximation of textual quality alignment. Let $\Delta q_{i}$ denote the expected quality improvement when offloaded to the cloud, and $(\Delta l_{i},\Delta k_{i})$ represent the additional latency and API cost, respectively. We define the normalized cost term as:
@@ Appendix D Supplementary Experiments
Effectiveness of SFT: As shown in Table  6, our SFT Planner achieves an accuracy of 22.00% on GPQA, outperforming the planner based on the Llama3.2-3B base model (20.00% accuracy). This demonstrates the effectiveness of our distillation and fine-tuning process in generating high-quality, logically sound plans.
Parallelization Advantage: At the same time, Table  6 shows the SFT Planner achieves a 34.3% compression ratio. The compression ratio $R_{\mathrm{comp}}$ is defined as follows:

### v4

@@ Abstract
Edge-cloud collaborative inference is crucial for LLM-powered edge devices, as on-device models often lack the required reasoning capability, while cloud-only inference can be costly and slow under strict latency and token/API budgets. However, existing edge-cloud collaboration methods typically route input tasks based on their estimated difficulty. These static, coarse heuristics overlook subtask dependencies, missing opportunities for parallel execution and budget-adaptive routing. To this end, we propose HybridFlow, a resource-adaptive edge-cloud inference framework that enables parallel execution of interdependent subtasks. Specifically, we build a dependency-aware DAG for each input task, facilitating concurrent execution of subtasks once their dependencies are resolved, thereby reducing end-to-end latency. Additionally, we propose a dynamic benefit–cost utility model, optimizing the trade-off between accuracy, token/API cost, and latency in real-time. This dynamic routing minimizes unnecessary cloud usage while preserving reasoning quality. Across GPQA, MMLU-Pro, AIME24, and LiveBench-Reasoning, HybridFlow improves the cost-accuracy trade-off, reducing latency and cloud API usage while maintaining competitive accuracy. Code: HybridFlow
@@ 2 Related Work
Algorithm 1 Dependency-Aware Subtask Scheduling with Budget-Adaptive Routing
Input: User query $Q$
@@
Stage 1: Task Decomposition and DAG construction
$(T,E)\leftarrow\mathrm{Decompose}(Q;M_{P})$ ;
@@
// fallback to chain if invalidInitialize $\deg(\cdot)$ as in-degree in $G$ Initialize frontier queue $\mathcal{Q}\leftarrow\{x\in T:\deg(x)=0\}$ Stage 2: Budget-Adaptive Routing and ExecutionInitialize $C_{\mathrm{used}}(t)\leftarrow 0$ while $\mathcal{Q}$ not empty do
   Pop a ready subtask $x$ from $\mathcal{Q}$. Obtain subtask embedding: $z(x)=\texttt{embedding}(x)$ Predict utility score $\hat{u}(x)=f_{\theta}(z(x),C_{\mathrm{used}}(t))$. Compute adaptive threshold $\tau_{t}$ from current resource usage. if $\hat{u}(x)>\tau_{t}$ then
    $R_{x}\leftarrow M_{\mathrm{cloud}}(x)$; ;
    // Cloud executionUpdate $C_{\mathrm{used}}(t)$.
@@
    $R_{x}\leftarrow M_{\mathrm{edge}}(x)$; ;
    // Edge executionUpdate $C_{\mathrm{used}}(t)$.
  Append $R_{x}$ to global context and push newly unblocked subtasks into $\mathcal{Q}$.
Stage 3: Final AggregationCombine all sub-results in topological order to form $R$. return $R$
@@ 3.1 Preliminaries
  • $\Delta k_{i}$: additional API usage cost (in price units) incurred by cloud execution.
In practice, $\Delta q_{i}$ is obtained via an offline outcome-based credit assignment procedure. For each subtask $i$, we construct paired executions that differ only in whether subtask $i$ is processed on the cloud or on the edge while keeping the rest of the pipeline unchanged.
@@
To place latency and API usage on a common scale, we define a normalized offloading cost. For each subtask $t_{i}$, the normalized cost of offloading is
@@
where $l_{\max}^{\mathrm{sub}}$ and $k_{\max}^{\mathrm{sub}}$ are per-subtask upper bounds on additional latency and API cost. Typically $\Delta l_{i}\leq l_{\max}^{\mathrm{sub}}$ and $\Delta k_{i}\leq k_{\max}^{\mathrm{sub}}$, so $c_{i}$ is bounded and comparable across subtasks. We use a binary variable $r_{i}\in\{0,1\}$ to indicate whether $t_{i}$ is offloaded to the cloud ($r_{i}=1$) or executed on the edge ($r_{i}=0$). Intuitively, subtasks with larger $\Delta q_{i}$ and smaller $c_{i}$ are more attractive to offload.
@@ 3.3 Utility-based Subtask Routing
$\hat{u}_{i}=\sigma\!\big(f_{\theta}(z_{i},C_{\mathrm{used}}(t))\big)\in(0,1),$ (8)
where $\sigma$ is the sigmoid function and $C_{\mathrm{used}}(t)=\sum_{j\leq t}r_{j}c_{j}$ is the cumulative normalized cost used so far. The estimate $\hat{u}_{i}$ is trained to approximate the ideal utility $u_{i}$ in Def. 3.2.
@@
where $\eta>0$ is a step size. When the system overspends, $\lambda_{t}$ increases and the routing policy becomes more conservative.
@@ 4.1 Experiment Setup
We evaluate HybridFlow on four reasoning benchmarks spanning mathematical, scientific, and general domains: GPQA (Rein et al., 2024); AIME24; MMLU-Pro (Wang et al., 2024b), and LiveBench-Reasoning (White et al., 2025). We select three metrics that jointly capture reasoning quality and system efficiency. (i) $Acc$ measures the correctness of reasoning by comparing model outputs with gold answers across all benchmarks. (ii) $C_{\mathrm{time}}$ denotes the end-to-end inference latency per query, including decomposition, routing, and execution. (iii) $C_{\mathrm{API}}$ quantifies the API cost consumed by cloud LLM calls, reflecting API efficiency and cost.
@@
HybridFlow deploys Llama3.2-3B on the edge device in two roles: (i) as the planner $M_{P}$ that decomposes each query into a task-level DAG, and (ii) as the edge executor $M_{\mathrm{edge}}$ that runs subtasks not offloaded. Each subtask is encoded with qwen3-embedding-0.6b into an embedding $z_{i}$, which is fed to a lightweight router: a two-hidden-layer MLP that predicts an offline utility estimate $\hat{u}_{i}$. The router is offline warm-started with AdamW (learning rate $1\times 10^{-4}$) by regressing to profiled utility targets. During inference, HybridFlow performs online adaptation with dual-threshold updates that track budget consumption; an online calibration head is updated from partial feedback via a contextual-bandit strategy. Routing decisions use the current threshold to assign each ready subtask to either $M_{\mathrm{edge}}$ or the cloud model $M_{\mathrm{cloud}}$ (GPT-4.1, via API). All edge-side computation (planning, local execution, and embedding) runs on a single NVIDIA RTX 3090 GPU. All LLM use a fixed temperature of $0.6$.
@@ 4.2 Results
While HybridFlow’s accuracy approaches the top-performing CoT with GPT-4.1, it does so with dramatically lower API consumption and latency. Table 2 shows that HybridFlow achieves the lowest average $C_{\mathrm{API}}$ (0.0088) among all collaborative baselines, indicating high API efficiency in cloud usage. When considering both accuracy and efficiency jointly, HybridFlow consistently dominates alternative hybrid systems: it improves accuracy relative to HybridLLM and DoT while simultaneously reducing latency and API cost. This demonstrates that HybridFlow effectively balances the strengths of edge and cloud models through adaptive routing and parallelized task execution, achieving a superior overall accuracy–efficiency frontier.
Table 3: Ablation of routing strategies on GPQA. We report the subtask offload rate, end-to-end accuracy, latency, cloud API cost $C_{\text{API}}$, normalized total cost $c$ (lower is better), and the unified utility $u$ (higher is better; see Sec. LABEL:sec:metric). HybridFlow-Chain disables DAG parallelism and executes subtasks sequentially, isolating the effect of routing from scheduling. HybridFlow achieves the best accuracy–cost trade-off, yielding the highest utility.
@@ 4.3 Ablation Study
Table 3 compares our adaptive router with naive and deterministic allocation strategies. Executing all subtasks on the edge model (Edge) eliminates cloud cost but attains only 25.54% accuracy, while full cloud execution (Cloud) achieves the highest accuracy (57.28%) at higher latency (18.26 s) and API cost ($C_{\text{API}}{=}0.0185$). The Random baseline reduces API cost ($C_{\text{API}}{=}0.0075$) but yields lower accuracy (46.00%), indicating inefficient use of cloud budget. A fixed-threshold policy improves accuracy to 51.62% with modest cost, yet remains inferior to our learned routing. HybridFlow-Chain further shows that routing alone is insufficient: disabling dependency-aware parallelism reduces performance (50.62% accuracy; utility 0.6095). In contrast, HybridFlow achieves 53.33% accuracy at the same low API cost as Random ($C_{\text{API}}{=}0.0075$) and attains the highest utility (0.7940), demonstrating a better accuracy–cost trade-off under our unified metric.
@@ Appendix A Notations
$C_{\mathrm{used}}(t)=\sum_{j\leq t}r_{j}c_{j}$ Cumulative normalized cost at time $t$.
@@ Appendix C Implementation Details
    (Well-formed outputs) At least one node is labeled Generate, and every Generate node has no outgoing edges (i.e., it is a sink in $G(Q)$). We require exactly one Generate sink node that produces the final answer.
@@
  6. 6.
    (Dependency consistency) For any edge $t_{j}\to t_{i}$, the output of $t_{j}$ is referenced by $t_{i}$. Concretely, each subtask $t_{i}$ declares a set of required symbols $\mathrm{Req}(t_{i})$ and produced symbols $\mathrm{Prod}(t_{i})$ (parsed from the XML plan); we require $\mathrm{Req}(t_{i})\subseteq\bigcup_{t_{j}\in P_{i}}\mathrm{Prod}(t_{j})$.
@@
Validation and repair.
After parsing the XML plan, we validate it using the rules in Definition C.2 (acyclicity, rootedness, reachability, well-formed outputs, size constraint) and an additional dependency-consistency check based on $\mathrm{Req}(\cdot)$$\mathrm{Prod}(\cdot)$ symbols. If validation fails, we apply a bounded, deterministic repair procedure: (i) remove ill-typed edges (violating dependency consistency), (ii) break cycles by removing the lowest-confidence edge [^1] We define confidence as the planner’s self-reported score per edge in the XML; when unavailable, we use a fixed priority order., (iii) enforce rootedness/reachability by attaching orphan nodes to the root, (iv) if the plan is still invalid after $R_{\max}$ iterations, we fall back to a chain plan (sequential execution). We use $R_{\max}=2$ in all experiments to bound overhead.
Table 5: Planner DAG validity and repair statistics across benchmarks. Valid denotes plans passing validation without repair; Repaired denotes plans fixed within $R_{\max}$ iterations; Fallback denotes sequential chain plans.
Benchmark Valid (%) Repaired (%) Fallback (%) #nodes (avg)
GPQA 76 14 10 4.47
LiveBench-Reasoning 78 13 9 4.32
Table 5 reports the planner’s decomposition reliability across benchmarks. We observe that the majority of queries yield a valid DAG without any repair (76–78%), indicating that the prompt-based planner typically produces well-formed, executable dependency structures under our constraints (Definition C.2). For an additional 13–14% of queries, the initial plan violates one or more validity checks (e.g., minor reachability or acyclicity issues) but can be deterministically repaired within $R_{\max}$ iterations; these repaired plans are subsequently executed as DAGs. Only 9–10% of queries trigger the fallback-to-chain mechanism after bounded repair attempts, which guarantees robustness while keeping the planner overhead controlled. Finally, among the executed DAG plans (Valid+Repaired), the average number of subtasks is approximately 4.3–4.5, suggesting that the planner produces moderately granular decompositions that expose parallelism without excessive fragmentation, consistent with our $n_{\max}=7$ cap.
@@
To train the proposed resource-aware router, we construct an offline profiling dataset from 2,0002{,}000 sampled queries drawn from two benchmarks: MMLU-Pro (different from the main test samples) and Math500 (covering general knowledge reasoning, targeting structured, multi-step reasoning). For each query, we perform the complete pipeline of task decomposition, subtask allocation, and execution using both the edge and cloud models. During this process, we record the response quality, inference latency, and API cost for each subtask.
For each subtask $i$, we construct paired executions that differ only in whether subtask $i$ is processed on the cloud or on the edge while keeping the rest of the pipeline unchanged. We then evaluate the two resulting end-to-end answers with a task-specific verifier (e.g., exact-match / multiple-choice correctness or a rubric-based judge) to compute the marginal correctness gain of cloud execution. Concretely, $\Delta q_{i}$ is defined as the difference between the final outcome scores of the two trajectories (cloud vs. edge for subtask $i$). This offline profiling is performed once on a held-out set and incurs no additional overhead during online inference.
In practice, each query is decomposed into 4455 subtasks on average (we constrain the planner to generate fewer than $7$ subtasks). For each query, we execute every subtask once on the edge model and once on the cloud model, and cache the resulting subtask outputs. Since the planner is fixed and does not re-plan conditioned on upstream outputs, we can efficiently construct mixed edge–cloud executions by recombining these cached outputs. We then sample a set of routing vectors and evaluate the final task outcome for each recombined execution. Finally, we estimate the contribution of each subtask by averaging its marginal effect over the sampled contexts, i.e., comparing outcomes when toggling that subtask between edge and cloud while keeping the remaining routing choices unchanged. Overall, this reuse-and-recombine strategy provides a reasonable approximation of subtask-level effects while substantially reducing profiling cost compared to rerunning the full pipeline for every counterfactual.
Let $\Delta q_{i}$ denote the expected quality improvement when offloaded to the cloud, and $(\Delta l_{i},\Delta k_{i})$ represent the additional latency and API cost, respectively. We define the normalized cost term as:
@@ Appendix D Supplementary Experiments
Effectiveness of SFT: As shown in Table  7, our SFT Planner achieves an accuracy of 22.00% on GPQA, outperforming the planner based on the Llama3.2-3B base model (20.00% accuracy). This demonstrates the effectiveness of our distillation and fine-tuning process in generating high-quality, logically sound plans.
Parallelization Advantage: At the same time, Table  7 shows the SFT Planner achieves a 34.3% compression ratio. The compression ratio $R_{\mathrm{comp}}$ is defined as follows: